|-6| + 3|4| + |-5| -0 . 6 =
Answer by HyperBrain(694)

Remember:
|-2|=2
|2|=2
|-3|=3
gets?

so,
|-6| + 3|4| + |-5| -0.6 =-6+3(4)+5-0.6=-6+12+5-0.6=6+5-0.6=11-0.6=10.4

Power up,
HyperBrain!